### v1

# Coherent synchrotron radiation by excitation of surface plasmon polariton on near-critical solid microtube surface

Abstract
Coherent synchrotron radiation (CSR) is vital for developing powerful ultrashort light sources. We introduce a CSR generation mechanism using surface plasmon polaritons (SPPs) resonantly excited on a solid, near-critical-density microtube. A high-intensity, circularly polarised laser pulse, propagating along the microtube axis, efficiently couples the cylindrical SPP modes. This process creates azimuthally structured, rotating electromagnetic fields. These rotating fields subsequently confine, modulate, and directly accelerate surface electrons, causing them to emit CSR in the Valilov-Cherenkov angle. We further demonstrate that by improving the azimuthal symmetry, the helical modulation enables CSR emission across all azimuthal directions, significantly enhancing radiation intensity even when full coherence is imperfect. The harmonics can be well isolated for a high charge beam. Our full 3D Particle-in-Cell simulations indicate this scheme can generate X-rays with coherence enhanced by up to two orders of magnitude compared to incoherent emission.
pacs:
Valid PACS appear here
Coherent synchrotron radiation (CSR) represents a unique regime of electromagnetic (EM) radiation emitted when relativistic charged particles—typically electrons—undergo accelerated motion in a collective, spatially ordered configuration [1]. Unlike conventional incoherent synchrotron radiation, where individual particles emit independently, CSR arises from the constructive interference of EM fields at specific wavelengths with fixed phase relationships between radiating electrons. The coherence results in a quadratic increase in power with the number of electrons, rather than the linear increase seen in incoherent cases. This property makes CSR highly advantageous for applications requiring high-power radiation sources in the terahertz (THz) and far-infrared regimes, such as ultrafast spectroscopy [2], particle acceleration diagnostics [3, 4], and next-generation light sources [5, 6].
Figure 1: Schematic: a CP laser pulse (yellow) enters the vacuum channel of the microtube (grey) and excites SPPs while being scattered at the sharp edge. The rotation mode $m=\pm 1$ (green-yellow) can efficiently couple with the laser field and accelerates the trapped electrons(sphere, blue-to-red colour presents the energy) into a spiral motion along the solid surface for the CSR generation (yellow-red cone).
The rapid progress in relativistic laser-plasma interactions has promoted the emergence of new compact radiation sources [7, 8, 9, 10]. Furthermore, an unexplored regime of superradiance has been recently discovered where coherent radiation can be obtained by specific modulation of the bunch to make the quasiparticle travel luminally or superluminally [11], which is distinct from the conventional superradiance mechanism [12]. This superluminal motion causes radiated phase fronts emitted at different times to intersect, forming an optical shock or caustic. At the crossing angle, the spherical waves emitted by other particles in the bunch have the same phase and thus interfere constructively. This new theory unlocks novel CSR in plasma-based accelerators [13].
A high-intensity laser pulse interacting with an interface between a metal and a dielectric medium (e.g., air or vacuum) can induce collective oscillations of free electrons. These oscillations, known as surface plasmons (SPs), are characterised by the plasma frequency, $\omega_{p}=\sqrt{4\pi n_{e}e^{2}/m_{e}}$, where $n_{e}$ is electron density, $e$ is electron charge, $m_{e}$ is electron rest mass [14]. The excitation of SPs has been recently proposed for generating TV/m-level wakefields in solid microtubes for relativistic particle acceleration [15]. When these SPs couple with incident EM waves, they form surface plasmon polaritons (SPPs)—hybrid light-electron modes that propagate along the interface, also termed surface plasmon waves [16]. Direct excitation of SPPs on a smooth, infinite interface is inherently inefficient due to the momentum mismatch between the incident light and the SPP modes [17]. However, in finite-sized solid microtargets, sharp edges can serve as effective scatterers, supplying the necessary momentum to bridge this gap and facilitate efficient SPP excitation [18]. The EM fields associated with SPPs can accelerate electrons along the target surface to relativistic energies. Recent studies have shown that irradiating the edge of a flat microtape with a high-intensity, linearly polarised laser pulse can generate strong SPP, producing high-energy electrons and high-flux X-ray radiation [19].
In this letter, we first show that SPPs can be directly excited at the cylindrical solid interface between plasma and vacuum by a circularly polarised (CP) laser pulse propagating paraxially through a microtube of near-critical density, as illustrated in Fig. 1. The cylindrical geometry enables the CP laser to couple efficiently with a rotating SPP mode whose phase velocity approaches the speed of light in vacuum. This resonant coupling provides $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level SPP fields for the modulated trapping, confinement and stable acceleration of relativistic electrons over several Rayleigh lengths. By matching the channel radius with the SPP field, the modulated high-energy electrons stably rotate near the surface and can produce CSR in the Valilov-Cherenkov angle. We further show that with an azimuthally symmetric distribution, the modulated beam can emit coherent radiation in all azimuthal directions. It also indicates that the harmonics can be well isolated. Full 3D particle-in-cell (PIC) simulations have been used to verify the fundamental process. Numerical simulations show significant coherence enhancement of photon emission in comparison with the incoherent case, with the radiation spectrum of critical energy at tens of $\mathrm{k}\mathrm{e}\mathrm{V}$.
A microtube can be made of nano-materials, e.g. longitudinally aligned carbon nanotube (CNT) forest, which can provide the solid plasma with flexible density in the range of $10^{19}-10^{24}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$ [20]. The SPP field is shown in Fig. 2, which is obtained from a PIC simulation where a microtube of radius $a=1.5\leavevmode\nobreak\ $\mu\mathrm{m}$$ is used. A $20\leavevmode\nobreak\ $\mu\mathrm{m}$$-long vacuum section is placed at the head of the target to initialise the laser pulse and provides a sharp annular edge. As the bulk size of the tube is much larger than the inner radius of a CNT, e.g. up to several $\mathrm{n}\mathrm{m}$, the effect from the inner structure of CNTs is neglected. A CP laser pulse of wavelength $\lambda_{l}=0.8\leavevmode\nobreak\ $\mu\mathrm{m}$$ and normalized strength $a_{0}=6$ propagates in the $z$ direction and is focused to the entrance of the microtube at $z=0\leavevmode\nobreak\ $\mu\mathrm{m}$$, with root-mean-square (RMS) waist $w_{0}=2.0\leavevmode\nobreak\ $\mu\mathrm{m}$$, and duration $\sigma_{\tau}=20.0\leavevmode\nobreak\ $\mathrm{f}\mathrm{s}$$. These parameters give a peak field of $24.0\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$, or $13.6\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$ on the inner surface. This is sufficient for the pulse front to instantaneously ionise electrons of the material within half a cycle. Therefore, the material properties, e.g., lattice structure or optical absorption, are unimportant for electron dynamics. For ultrashort laser pulses, e.g. duration $L\ll 100\leavevmode\nobreak\ $\mathrm{f}\mathrm{s}$$, the plasma expansion can be neglected and the solid target can be assumed to maintain the sharp-edge cold plasma profile. These conditions allow the propagation of SPPs on the vacuum-plasma interface.
The simulations are carried out with the code WarpX [21]. The dimensions of the moving window are $8\leavevmode\nobreak\ $\mu\mathrm{m}$\times 8\leavevmode\nobreak\ $\mu\mathrm{m}$\times 30\leavevmode\nobreak\ $\mu\mathrm{m}$$, comprising $512\times 512\times 1024$ cells in the $x$, $y$, and $z$ directions, respectively. Each cell contains 64 macro particles, which are sufficient to solve the electron dynamics effectively. The microtube is initially modelled as a uniform distribution of neutral carbon atoms, with their ionisation energies adjusted to reflect CNT properties—specifically the weak $\pi$-bond and $\sigma$ (C–C) bond energies. The initial carbon density is $n_{C}=2\times 10^{21}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$ which gives $\omega_{p}$ close to the laser frequency $\omega_{l}$. The field ionisation is implemented using the Ammosov-Delone-Krainov (ADK) method  [22].
Figure 2: (a) Electron density distribution $n_{e}$ (black colour) and $E_{z}$ (blue-red colourmap) in $yz$-plane sliced at $y=0$ while the laser pulse entering the microtube, respectively. (b) Snapshots of $E_{z}$ in $xy$-plane sliced at $z=5\leavevmode\nobreak\ $\mu\mathrm{m}$$ corresponding to (a). (c)-(e) Snapshots of $E_{z}$, $E_{r}$ and $E_{\phi}$ in $xy$-plane at the same phase after the laser pulse propagating $z=41\leavevmode\nobreak\ $\mu\mathrm{m}$$ in the tube. The solid black and dashed red lines in (e) present the line plots of $E_{z}$ along $y=0\leavevmode\nobreak\ $\mu\mathrm{m}$$ from PIC and analytical solutions, respectively.
The vertical solid surface at $z=0\leavevmode\nobreak\ $\mu\mathrm{m}$$ scatters the incident CP laser pulse and enables the direct matching with the SPP mode if the resonant condition $k_{l}=k_{sp}$ are satisfied, where $k_{l}$ and $k_{sp}$ are wave numbers of laser and SPP field, respectively. This condition is efficient near critical density, as shown in Fig. 2(a) and (b). The excited SPP fields can stably propagate with ultrahigh strength in $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level, e.g. $4.5\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$ in our simulations as shown in Fig. 2 (c)-(e) where the components are sliced after the laser pulse propagating $z=41\leavevmode\nobreak\ $\mu\mathrm{m}$$ in the microtube. Some electrons can be trapped into the SPP field and accelerated to relativistic speed by the longitudinal component $E_{z}$ and, at the same time, confined in a small vicinity near the surface by the transverse component $E_{r}$ if they have sufficient energy, as shown in Fig. 2 (c). Due to the rotating phase front of the CP laser pulse, the electron beam is modulated to be helical during the trapping process, as shown in Fig. 3(a). The azimuthal component $E_{\phi}$ drives the modulated beam to rotate.
The PIC results can be confirmed theoretically by analysing the SPP excitation in cylindrical geometry, $(r,\phi,z)$ [23]. In general, a pure transverse magnetic (TM) mode (e.g. $B_{spp,z}=0$ and $E_{spp,z}\neq 0$) of SPP field can exist on the cylindrical surface. The $z$-component of a SPP mode can be given by solving Maxwell equations, as $E_{spp,z}^{(m)}(r,\phi,z)=Af_{m}(\kappa r)e^{i(k_{z}z+m\phi-\omega t)}$, where $A$ is a constant representing the field amplitude. $m$ is the SPP mode number and $\kappa=\sqrt{k_{z}^{2}-\epsilon_{d}\omega^{2}/c^{2}}$ with $k_{z}$ the wavevector along the tube axis, $\omega$ the frequency, $c$ is the speed of light in vacuum and $\epsilon_{d}$ permittivity of the medium. $f_{m}(\kappa r)$ presents the radial field profile and depends on the medium properties. In the vacuum channel, $r<a$, the field is presented by the modified Bessel functions of the first kind as $f_{m}(\kappa r)=I_{m}(\kappa_{1}r)$ since the field is finite as $r\to 0$ and $\epsilon_{d}=\epsilon_{1}=1$. In the plasma, $r>a$, the field is presented by the the modified Bessel functions of the second kind as $f_{m}(\kappa r)=K_{m}(\kappa r)$ the field decays to zero as $r\to\infty$ and $\epsilon_{d}=\epsilon_{2}=1-n_{e}/n_{c}(\omega)$ where $n_{c}$ is the critical density for the frequency $\omega$ as $n_{c}(\omega)=m_{e}\omega^{2}/4\pi e^{2}$. PIC simulation agrees well with the analytical solution as shown in Fig. 2(c)-(e). The other components can be obtained from the axial components using Maxwell’s equations, for example, the electric components as $E_{ssp,r}^{(m)}=(ik_{z}/\kappa^{2})\partial E_{spp,z}^{(m)}/\partial r$ and $E_{spp,\phi}^{(m)}=imk_{z}E_{spp,z}^{(m)}/\kappa^{2}r$ for different regions, respectively. See more details in the Supplementary Material.
The dispersion relation can be derived by considering the continuity of the SPP field at the interface of plasma and vacuum as
$\frac{I_{m}^{\prime}(\kappa_{1}a)}{\kappa_{1}I_{m}(\kappa_{1}a)}=\frac{K_{m}^{\prime}(\kappa_{2}a)}{\kappa_{2}K_{m}(\kappa_{2}a)}\mathrm{,}$ (1)
where $\kappa_{1}=\sqrt{k_{z}^{2}-\omega^{2}/c^{2}}$ and $\kappa_{2}=\sqrt{k_{z}^{2}-\epsilon_{2}\omega^{2}/c^{2}}$. For a large $a$, the dispersion relation in Eq. (1) resembles for $m=0$ as $k=(\omega/c)\sqrt{\epsilon_{2}/(1+\epsilon_{2})}$, which is the dispersion relation of SPP on the flat surface [14]. In the $m=0$ mode, the phase velocity $v_{ph}$ evolves similarly with that in planar geometry and only approaches $c$ as $n_{e}$ is much higher than $n_{c}$. However, in $m=1$ mode, $v_{ph}$ is very close to $c$ even for a low plasma density. This feature enables the $m=1$ mode to be used for relativistic electron acceleration in a near-critical plasma.
Figure 3: PIC results: 3D (a) and front (b) views of the high-energy ($\mathcal{E}>88\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$) electron beam modulation after the laser propagates $z=40\leavevmode\nobreak\ $\mu\mathrm{m}$$ in the microtube. The arrows in (b) present the transverse momentum field ($u_{x},u_{y}$). (c): Energy evolution of the high-energy electrons in the beam (grey lines) and the mean value (black line) as a function of propagation distance. (d) Trajectories of these electrons. The red line represents one of the electrons in the beam.
The constant $A$ is proportional to the laser field amplitude $E_{0}$ and the overlap efficiency factor $\mathcal{O}(a,w_{0},\kappa,m=1)$ as $A\approx\mathcal{O}\cdot E_{0}$. From the PIC simulations as shown in Fig. 2, $\mathcal{O}\simeq 0.19$, which implies that $100\leavevmode\nobreak\ $\mathrm{T}\mathrm{W}$$-level laser pulse can drive $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level SPP field. The efficiency is orders of magnitude higher than that with a linearly polarised pulse where $\mathcal{O}\sim 10^{-3}$ [15]. The absolute value of $A$ depends on the incident laser power and the overlap integral between the incident laser field profile at the coupling region and the SPP mode profile, as
$A\propto\int_{0}^{a}\int_{0}^{2\pi}\left(\mathbf{E}_{spp,\perp}\cdot\mathbf{E}_{laser,\perp}^{*}\right)\bigg{|}_{z=0}r\,dr\,d\phi\mathrm{,}$ (2)
where $\perp$ denotes components transverse to the z-axis. Eq. (2) indicates that, except $m=0$ mode, the SPP field with the azimuthal phase dependence $e^{im\phi}$ allows the enhanced coupling if the incident field matches the mode. As a result, the efficient SPP excitation depends on how the laser field matches the azimuthal symmetry of the SPP mode. The coupling efficiency can be determined by $(\mathbf{E}_{spp}^{(m)}\propto e^{im\phi})\cdot(\mathbf{E}_{\text{laser}}^{*}\propto e^{-il\phi})\sim e^{im\phi}\times e^{-il\phi}=e^{i(m-l)\phi}$ where $l$ is the azimuthal index of the laser field. For a CP laser pulse, $l=\pm 1$ and the electric field $\mathbf{E}_{\text{laser}}(r,\phi,z,t)\approx E_{0}(\hat{\mathbf{r}}+i\hat{\boldsymbol{\phi}})g(r)e^{\pm i\phi}e^{i(k_{l}z-\omega_{l}t)}$ where $g(r)$ represents the spatial envelope. The symbol $\pm$ denotes the right (RHC) or left (LHC) hand circular polarisation, respectively. The azimuthal structure of RHC or LHC polarisation can match the azimuthal phase dependence of rotation SPP mode $m=1$ or $m=-1$, as $e^{\pm i\phi}\times e^{\mp i\phi}=1$, respectively. This leads to strongly enhanced coupling of these modes as seen in Fig. 2. For the other modes, e.g. $m\neq 1$, integrating $e^{i(m\pm 1)\phi}$ over $2\pi$ gives zero, and the excitation is not efficient.
The resonant excitation of SPP leads to highly efficient electron acceleration and modulation as shown in Fig. 3(c) where the energy gain is $4.0\leavevmode\nobreak\ $\mathrm{T}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{m}$$ in peak and the acceleration can be maintained for $40\leavevmode\nobreak\ $\mu\mathrm{m}$$ or 3 Rayleigh length $Z_{R}$. Due to the continuous trapping process, the energy spectrum is broad. Part of electrons, e.g. $18.7\leavevmode\nobreak\ $\mathrm{p}\mathrm{C}$$ or about $10\%$ of the total trapped electrons, can be accelerated to high energy $>90\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$ due to the early trapping. These electrons can rotate around the channel axis, being confined closely along the interface with the helical modulation as shown in Fig. 3(a) and (b). The modulation is introduced during the injection process by the laser field with the frequency $\omega_{m}=\omega_{l}$. The rotation is driven by the azimuthal components, and the frequency $\omega_{r}$ is estimated as $\omega_{r}=\sqrt{e(E_{\phi}+\beta_{z}cB_{r})/m_{e}a}$. With the value of fields from PIC simulations, $\omega_{r}\simeq 2\pi\times 10^{13}\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$ or $\lambda_{r}=2\pi c/\omega_{r}\simeq 30\leavevmode\nobreak\ $\mu\mathrm{m}$$, as shown in Fig. 3(d). The normalised phase velocity of the modulation is given as $\beta_{ph}\simeq 1/(1-\omega_{r}/\omega_{m})>1$, which allows the constructive interference.
Figure 4: (a) Radiation spectrum coherently calculated by the trajectories of the electrons from a PIC simulation in Fig. 3. Insert (b): Line plot along Valilov-Cherenkov angle, $\varphi_{vc}\simeq 0.23$, indicated by horizontal yellow dashed line. The dashed white curves are theoretical resonant frequencies of different harmonics.
The helically modulated oscillation of the electron beam can result in the CSR in the form of optical shock along the Valilov-Cherenkov angle, $\cos\varphi_{vc}\simeq(1-\omega_{r}/\omega_{m})$ [11]. The radiation spectral lines are centred at frequencies $\tilde{\omega}_{n}=(n\pm 1)\omega_{r}/(1-\beta_{z}\cos\theta)$ where $n$ is an integer representing the harmonic number and $\beta_{z}$ the normalised longitudinal velocity of the beam. $\theta$ the angle with $z$-axis in $yz$-plane. For $\beta_{z}\sim 1$, $\tilde{\omega}_{n}(n=0)\simeq\omega_{m}$ along in Valilov-Cherenkov angle $\theta=\varphi_{vc}$. The radiation intensity is proportional to the square of the number of electrons inside the beam, $N$. With the parameters used in this letter, $\varphi_{vc}\simeq 0.23$. The radiation spectrum is shown in Fig. 4 (a) and (b), which is coherently calculated by using the wave optical method [24] with a virtual spherical detector in the far field where the radiation spikes appear in $\varphi_{vc}$ angle at the harmonic frequencies $\tilde{\omega}_{n}$. However, there are no advantages in modulating the bunch at the radiation frequency.
Now, we theoretically consider a specific modulation, azimuthally symmetric modulation, where the initial phase of the electron trajectories is azimuthally uniform. The trajectory of an electron (denoted by $j$) inside the electron beam can be written as $\bm{r}_{j}=(x_{j},y_{j},z_{j})=a\left(\bm{\epsilon}e^{i\omega_{r}(t-t_{0j})+i\psi_{j}}+c.c.\right)/2+\bm{\eta}\beta_{z,j}c(t-t_{0j})$, where the polarization vectors are defined as $\bm{\epsilon}=(1,\pm i,0)$ and $\bm{\eta}=(0,0,1)$. $t$ is the retarded time and $t_{0j}$ the initial injection time. In general, the initial phase $\psi_{j}$ depends on the injection time as $\psi_{j}=\omega_{m}t_{0j}$, which introduces the helical modulation with azimuthal symmetry. $\omega_{m}$ is the modulation frequency and $\omega_{m}=\omega_{l}$ in our case. Now, we further consider an additional condition, $\psi_{j}=\omega_{m}t_{0j}=2\pi j/N$, which implies that the electrons inside the beam are injected uniformly in the azimuthal phase. Note that this condition does not require the electrons to be close, as is the case with conventional superradiance [12]. Since the beam is ultra-relativistic, $\beta_{z,j}\simeq\beta_{z}\simeq 1$.
[…]
$\mathcal{M}=\begin{pmatrix}-(\sin^{2}\theta\sin^{2}\phi+\cos^{2}\theta)&\sin^{2}\theta\cos\phi\sin\phi&\sin\theta\cos\theta\cos\phi\\ \sin^{2}\theta\cos\phi\sin\phi&-(\sin^{2}\theta\cos^{2}\phi+\cos^{2}\theta)&\sin\theta\cos\theta\sin\phi\\ \sin\theta\cos\theta\cos\phi&\sin\theta\cos\theta\sin\phi&-\sin^{2}\theta\end{pmatrix}\mathrm{.}$
Figure 5: (a) Theoretical radiation spectrum calculated coherently (upper) and incoherently (bottom) by using an electron beam of $N=100$ electrons uniformly distributed over six modulation periods, or $4.8\leavevmode\nobreak\ $\mu\mathrm{m}$$ long. The other beam parameters are the same as those from PIC simulations, as shown in Fig. 3. Insert (b) Line plot of (a) along axis $\theta=0$ and $\phi=\pi/2$. (c) Line plot of onaxis ($\theta=0$ and $\phi=\pi/2$) spectrum in Fig. 4(a).
Using the modulated electron trajectories, the vector field of radiation can be given by
$\begin{split}\bm{\mathcal{A}}&=\frac{NN_{r}\pi}{\omega_{r}}\mathcal{M}\cdot\sum_{n=-\infty}^{\infty}i^{n}\bigg{[}k_{r}a\bigg{(}\bm{\epsilon}J_{nN-1}(\zeta)\operatorname{sinc}(\omega_{-})+\\ &\bm{\epsilon^{*}}J_{nN+1}(\zeta)\operatorname{sinc}(\omega_{+})\bigg{)}+2\bm{\eta}\beta_{z}J_{nN}(\zeta)\operatorname{sinc}(\omega_{z})\bigg{]}\end{split}$ (3)
with $\omega_{\pm}=N_{r}\pi(\omega_{nN}\pm\omega_{r})/\omega_{r}$, $\omega_{z}=N_{r}\pi\tilde{\omega}_{n}/\omega_{r}$ and $\omega_{nN}=\omega(1-\beta_{z,j}\sin\phi\cos\theta)+nN\omega_{r}$. $\zeta=(\tilde{\omega}a/c)(\sin\theta\sin\phi-\sin\theta\cos\phi)$. See more details in the Supplementary Material. It is seen from Eq. (3) that the radiation spectral lines are centred at frequencies $\tilde{\omega}_{n}=(nN\pm 1)\omega_{r}/(1-\beta_{z}\cos\theta\sin\phi)$, which presents the CSR in all azimuthal directions rather than being confined to the angle $\varphi_{vc}$. It also indicates that the harmonic numbers are given by $nN$. This implies the harmonics can be well isolated. For sufficiently large $N$, such as a $\mathrm{p}\mathrm{C}$-level beam, only the fundamental harmonic ($n=0$) is present in the spectrum, as shown in Fig. 5(a). This characteristic offers a significant advantage in that the precise control over the radiation spectrum can be significantly enhanced. The on-axis radiation cut-off frequency is $\tilde{\omega}_{c}\simeq 3K_{0}\gamma^{2}\omega_{r}/2$ [25]. For an average $\gamma=150$, this yields $\hbar\tilde{\omega}_{c}\simeq 31.3\leavevmode\nobreak\ $\mathrm{k}\mathrm{e}\mathrm{V}$$, which aligns with the numerical simulations presented in Fig. 5 (a) and (b).
In realistic beams, such as those obtained from PIC simulations, modulated electrons are distributed over a finite radial extent of $\Delta r\simeq 0.15\leavevmode\nobreak\ $\mu\mathrm{m}$$, as shown in Fig. 3(b). This spatial spread disrupts the azimuthal symmetry, thereby impeding fully coherent radiation emission. Nonetheless, partial coherence can still be achieved. This occurs when a subset of electrons satisfies the necessary symmetry requirements, leading to an enhanced radiation intensity, as shown in Fig. 5(c), where the photon emission per electron is increased by almost two orders of magnitude compared to that predicted by incoherent calculations.
In summary, the proposed microtube can be fabricated from crystals that provide solid plasmas. A promising candidate is a nanostructured carbon nanotube forest, which offers tunable density and excellent plasma uniformity. The edge-coupling mechanism presented in this letter also enables the excitation of other modes by matching the azimuthal mode number $m$, for example, by a Laguerre-Gaussian laser [26]. The helical modulation of the electron beam can result in CSR, although it requires modulation of the fundamental radiation frequency. If the electron distribution is azimuthally uniform, the CSR can be generated in all directions. However, further investigation is required into optimising SPP excitation and acceleration to improve the azimuthal symmetry of the accelerated beam.
Acknowledgements.
Javier Resta-López acknowledges support by the Generalitat Valenciana under grant agreement CIDEGENT/2019/058. This work made use of the facilities of the N8 Centre of Excellence in Computationally Intensive Research (N8 CIR) provided and funded by the N8 research partnership and EPSRC (Grant No. EP/T022167/1). The Centre is coordinated by the Universities of Durham, Manchester and York.